OpenStudy (anonymous):

determine whether each function is linear or quadratic. identify the quadratic linear, and constant terms y=x+4 f(x)=x2-7 g(x)=-7(x-4) y=x(1-x)-(1-x2)

OpenStudy (anonymous):

please really need help

OpenStudy (anonymous):

y=x+4 and g(x) and y=x(1-x)-(1-x^2)[=x-x^2-1+x^2=x-1] are linear f(x) is quadratic

OpenStudy (anonymous):

the numbers are the constant terms, the x is the linear term and the x^2 is the quadratic term